# Relative equilibria for Scaling Symmetries and Central Configurations

### v1

@@ 7 Relative Equilibria of Scaling Symmetry for Cotangent Bundles
The action of the group $G=\mathbb{R}^{+}$ on $Q$ is given by $\Psi(g,\mathbf{q})=(g,g\mathbf{q})$. The tangent lifted action on $T^{\ast}Q$ is $\Psi^{\ast}(g,(\mathbf{q},\mathbf{\dot{q}}))=(g,(g\mathbf{q},g\mathbf{\dot{q}}))$, and the scaled cotangent lifted action is $\hat{\Psi}(g,(\mathbf{q},\mathbf{p}))=(g,(g\mathbf{q},g^{-1/2}\mathbf{p}))$. It is straightforward to verify that $\hat{\Psi}_{g}^{\ast}H=g^{-1}H$ and $\hat{\Psi}_{g}^{\ast}\omega_{0}=g^{1/2}\omega_{0}$, confirming that $\hat{\Psi}$ is a scaling symmetry.
@@
If $c=\frac{1}{2}$, the equation simplifies to
@@ Acknowledgments
We thank Shengda Hu and Cristina Stoica for useful discussions and remarks concerning this paper.

### v2

@@ 7 Relative Equilibria of Scaling Symmetry for Cotangent Bundles
The action of the multiplicative group $\mathbb{R}^{+}$ on $Q$ is given by $\Psi(g,\mathbf{q})=(g,g\mathbf{q})$. The cotangent lifted action on $T^{\ast}Q$ is $\hat{\Psi}(g,(\mathbf{q},\mathbf{{p}}))=(g\mathbf{q},g^{-1}\mathbf{p})$. Hence, $(\hat{\Psi}_{g})^{\ast}K(\mathbf{q},\mathbf{p})=g^{-2}K(\mathbf{q},\mathbf{p})$, and $\hat{\Psi}_{g}^{\ast}U(\mathbf{q})=\Psi_{g}^{\ast}U(\mathbf{q})=g^{-1}U(\mathbf{q})$. By Proposition 7.1 with $b=-1$ and $a=-2$ we take $c=(b-a)/2=1/2$, and hence the scaled cotangent lifted action is
$\hat{\Psi}^{c}(g,(\mathbf{q},\mathbf{p}))=(g\mathbf{q},g^{1/2}g^{-1}\mathbf{p})=(g\mathbf{q},g^{-1/2}\mathbf{p}).$
It is straightforward to verify that $(\hat{\Psi}_{g}^{c})^{\ast}H=g^{-1}H$ and $(\hat{\Psi}_{g}^{c})^{\ast}\omega_{0}=g^{1/2}\omega_{0}$, confirming that $\hat{\Psi}^{c}$ is a scaling symmetry.
@@
Taking $c=\frac{1}{2}$, the equation simplifies to
@@ Acknowledgments
We thank Shengda Hu, Corey Shanbrom, David Sloane and Cristina Stoica for useful discussions and remarks concerning this paper.